All systems are operational

Past Incidents

5th September 2023

No incidents reported

4th September 2023

No incidents reported

3rd September 2023

North Lincolnshire Barton-upon-Humber & Barrow-upon-Humber - Hard Down

We are currently investigating reports of lost connection in the Barton and Barrow areas of North Lincolnshire. Further updates will follow shortly. Please do not press the reset button on your router.

  • Resolved

  • The risk of hard down is now lowered. Continued on which remains open at this time.

  • Services are now restoring following the outage. This is being monitored and is considered at risk at the moment. Please refrain from completing speed tests whilst the connection re-establishes.

  • We are also receiving reports that connections in the DN39 are also affected. This is being investigated and further updates will follow shortly.

  • 2nd September 2023

    No incidents reported

    1st September 2023

    No incidents reported

    31st August 2023

    Lincolnshire North Thoresby Power Cut

    Power Cut in North Thoresby and Battery Back Up has unexpectedly saturated. Generator is being taken to site to resolve internet services for those whose homes have not lost power

  • Services have now restored. We will monitor services for a short time.

  • Power was believed to be affecting the whole village. As we are now aware this is not the case all resources are on route to the affected area with generators ETA to site 45 minutes. Further updates will follow

  • 30th August 2023

    No incidents reported

    29th August 2023

    Fibre fault and Scheduled Maintenance Liverpool

    Planned Start Date: 29/08/2023 Planned Start Time: 20:00 Planned End Date: 30/08/2023 Planned End Time: 06:00

    This window is classified as Affected.

    This means that service is expected to be hard-down between the hours of 20:00 to 06:00 and full service impacting.

  • From Vendor: 3rd party provider network fibres respliced restoring services. From ITY - Services seen to restore from 4.15am. Have been monitoring since and confirm services restored. Case can be closed and mark resolved pending RFO Report. Expected 5-7 working days.

  • Vendor Update:

    Dear Customer, We have just been advised that engineers are on site and are setting up. Further updates will be shared as we receive them. Regards.

  • Vendor Update:

    Dear Customer, Our 3rd party has provided an eta of 02:45 BST (GMT +1) for their engineers who have been instructed not to leave the site until we have confirmed that services have been restored. Further updates to follow. Regards.

  • Vendor Update -

    Our 3rd party has identified a work order that took place at a location that lines up with the fault seen and the timing of our services going hard down.

    The contractor that performed the work has all the records required to resolve the issue and has been escalated to dispatch engineers to resplice.

    The next update will be when we have been provided with an ETA.

  • Vendor Update:

    A conference call with NEOS senior management and fibre engineers along with our 3rd party provider is being arranged to formulate an action plan and a way forward to resolve this fault.

    A further update will be provided once this call has taken place.

  • Vendor Update: Dear Customer,

    We are still actively chasing our third-party vendor for an ETA for engineer to site and are pushing them through our internal escalation process.

    This fault is escalated internally to director level at NEOS.

    Our own fibre engineers remain on standby for a joint meet to assist with resolution.

    The next update will be in 1 hour unless we receive an significant update before then.

  • Latest Update:

    We are still actively chasing our third-party vendor for an ETA for engineer to site.

    As soon as this has been provided, we will dispatch our own fibre engineers for a joint meet to assist with resolution.

    The next update will be in 1 hour unless we receive an significant update before then.

  • Our vendor update:

    Dear Customer,

    Please note, we have now escalated this to the next level with our third-party vendor at level 5.

    Our NOC Management have been on a conference call with our vendor to discuss all previous findings to assist with resolution.

    We are still actively pushing them to provide an ETA for engineer on site and hope to have this soon.

    The next update will be in 1 hour.

  • Our Vendor Update:

    Dear Customer, Following our escalation with our third-party, they have now accepted our diagnostics and are sending out resources to our identified break location. As soon as we have engineers ETA we will inform you. Please be aware this has been escalated to our CEO and COO who are actively engaged.

    The next update will be in 1 hour.

  • The issue has been pinpointed to 3 fibres provided via a supplier other than VirgiN Media. Virgin Media have now cleared their major incident. A secondary provider has now been actioned to attend. They will accept all previous reports and testing information as their own and have not reset the service level agreement window of repair or resolution, so this is still being treated as a fibre fault that is 27 hours in. This is therefore at the highest escalation level and will be restored at the earliest opportunity. We do not have an estimated resolution time at this moment, but we do expect to have more quality information and detail within the next 2 hours.

  • Third Party Provider Update: We do apologise for the delayed update. We are actively escalating with our third party vendor for a plan of action on resolution. As soon as they have provided us with an ETR we will update you.

  • Third Party Provider Update: Please be advised that we are continuing to work alongside our 3rd party vendor to help them to establish the fault on their network and assist in any way possible to aid and speed up the resolution.

    Whilst the 3rd party are trying to identify their network issue, we do not currently have an ETR however, we are pushing for further updates as soon as possible.

    The next update will be at approximately 16:00 unless we have any updates from the vendor prior.

  • Third Party Provider Update: Testing has confirmed the break location that was identified last night was incorrectly associated with our network due to a crossed parallel fibre. All this testing information has now been shared with our 3rd party to speed up their investigation and fault resolution.

    Our engineers are still onsite and available for any co-op required with the 3rd party vendor where required.

    Next update will be at approximately 15:00 unless we have any further major developments before.

  • Third Party Provider Update: Further testing has identified their may be a further issue at an alternative location on leased fibres. We have now engaged the 3rd party vendor to investigate this issue and are awaiting their initial response.

    Our engineers are also continuing to work alongside Fibre teams to find a resolution as soon as possible including looking at the prospect of a re-route.

    Next update will be at 14:00 unless there are any other major developments before.

  • Third Party Provider Update: Our engineers and Tier 3 Fibre teams are continuing to investigate and work to find a resolution as soon as possible however, currently we have no further updates to share.

    We will continue to keep you updated as we receive any further details.

    The next update will be at approximately 13:00 unless we have any major changes before.

  • Third Party Provider Update: Currently our engineers are onsite working in conjunction with our Tier 3 Fibre Transmission team to investigate and find a resolution to this fault following failure of the initial repair.

    They are looking at several multiple options including alternative routes, additional working spares and potentially an installation of a strap cable to bypass the damage. However, due to the complex nature they are having to take all care when deciding on a solution to ensure further disruption is not caused.

    The next update will be at approximately 12:00 unless we have any further major developments.

  • Third Party Update: Engineers are continuing to investigate for a fix following the initial spare fibres not providing resolution.

    Due to the complex nature of the joints involved, engineers are having to take extra care in fibre identification.

    We will continue to update you on an hourly basis with the next update at approximately 11:00.

  • Third Party Provider Update: We have now had to engage further engineers to site due to regulations on working times with the current engineers. Full handover has been completed.

    Initial identification of spare fibres has been unsuccessful in resolving the issue, we are working to identify alternative options and a resolution as soon as possible

    Next update will be at approximately 10:00am.

  • Following completion of one joint being spliced, engineers are continuing further investigations at another joint trying to identify the spare fibres.

    We will continue to monitor for further updates from the engineers and be in touch when anything else is received.

    Next update will be at approximately 09:00.

  • Third party update -

    Engineers have tested and identified fibres which go the full distance and are currently prepping the fibres before splicing. Engineers have finished splicing at one joint and are moving to the next to complete their splicing.

  • Third Party Update -

    Our engineers have discovered that civil engineering works undertaken yesterday in Elgar Road Reading have damaged a 216f cable affecting customer services.

    Engineers have been working through the night to identify spare fibres but all the fibres tested so far have been damaged.

    Testing is ongoing with no ETR.

    Internetty -

    This now returns to fault status

  • Fibre engineers have entered a live fibre joint and are currently carrying out OTDR testing to pinpoint the location of the damage fibres. Due to working in the proximity of fibres carrying live traffic, this work is very delicate and can't be rushed. Next update at 05:00 unless significant progress has been made.

  • Both NEOS and 3rd party fibre engineers are at the location of the suspected fault and are walking the route to see if there are any obvious signs of civil engineering/building works taking place in the area. Please be advised that the SLA for this fault is P1 fibre with a fix time of upto 16 hours. Next update at 01:00

    We have already responded and called that 16 hours from point of first issue woudl extend beyond 6am and that this is not acceptable. They have not yet responded to our request for information.

  • Virgin Media engineers have OTDR tested the fibre and found the issue to be 800m from Reading Easynet PoP site, the 3rd party fibre provider have been requested to attend to pin point the issue.

    A further status update will be provided at 23:00

  • We can now confirm engineers have arrived on site and are in the process of locating the affected fibre. Once located, they will proceed to conduct the ODTR testing.

  • Statement from Third Party-

    We can now confirm that both engineers have arrived on site and are in the process of locating the affected fibre. Once located, they will proceed to conduct the ODTR testing.

    A further status update will be provided within 1 hour.

    We are now within Maintenance Window ITY2309002 but will continue to update you on the fibre break.

  • Engineers have yet to confirm arrival to site in order to begin ODTR testing - We have chased our supplier for more information on this. This fault will now extend into the maintenance window. We await further details on the estimated time for restoration of services under the maintenance window.

  • Although it has not yet been confirmed, we believe this outage will roll into the scheduled maintenance window due to commence 20:00hrs 29th August 2023. We will await confirmation from the third-party provider,

  • Statement from Neos Networks:

    Please be advised this fault is related to a wider outage that we believe is a suspected fibre break between Reading and Froxfield.

    We have engaged our 3rd party vendor regarding this as well as our own Fibre engineers to investigate further and to pin point the fault.

    There is currently no ETA however, we will endeavour to keep you updated as soon as we receive further information.

    Internetty -

    We are constantly requesting updates and understand customer frustrations. We are also looking into available alternatives to restore service faster if possible.

  • Urgent support request Created: 29/08/2023 15:15:14 BST

  • Priority 1 - Critical - has been raised with provider

  • If you are currently without service please leave any equipment on to ensure a smooth restoration, further updates/advice to follow. DO NOT FACTORY RESET ROUTERS, THIS WILL LEAD TO FURTHER DOWNTIME

  • Telehouse North 2, Coriander Avenue, London,E14 2AA to Liverpool Exchange BT Liverpool Royal,Royal House, 100 Wood Street, Liverpool, MERSEYSIDE,,L1 4LA - NEOS Networks reporting Hard down Fault. Awaiting their reponse